OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2).

OpenStudy (anonymous):

First find the slope of line y=3x+2. The line parallel to it has same slope then use y-y1=m(x-x1)

OpenStudy (anonymous):

2/3x-3?

OpenStudy (anonymous):

nope the slope y=3x+2 is M=3 THE EQUATION OF NEW LINE IS y-2=3(x-5), which is y-2=3x-15

OpenStudy (anonymous):

y=3x-13